OpenStudy (anonymous):

write an equation in slope-intercept form for the line with the given condition that contains the point (6,-2) Perpendicular to the line y=-3x+4

OpenStudy (anonymous):

Similar to your last question but since the lines are perpendicular the slope will be 1/3 because the slope of two perpendicular lines when multiplied togther equals -1.

OpenStudy (anonymous):

what is the equation?

OpenStudy (anonymous):

??????

OpenStudy (anonymous):

Oh OK... It is y+2 = 1/3(x - 6)

OpenStudy (anonymous):

Solve for y

OpenStudy (anonymous):

y can you solve, lol i don't understand this

OpenStudy (anonymous):

i'll give you a medal

OpenStudy (anonymous):

Clearly...I'll try to teach you. The slope intercept form is y - y1 = m(x - x1) So you know the slope is 1/3 and you know the coordinates of one point the line goes through. so you substitute the values into x1 and y1

OpenStudy (anonymous):

So now that you have the following equation y + 2 = 1/3(x - 6) y + 2 = 1/3x - 2 y = 1/3x - 4

OpenStudy (anonymous):

Do you sort of understand now?

OpenStudy (anonymous):

kinda its a little more clear

OpenStudy (anonymous):

thankyou for your help

OpenStudy (anonymous):

you are welcome